OpenStudy (anonymous):

We have: \[\int_0^1 \int_{3y^2}^{4-y^2}\int_{-\sqrt{z-3y^2}}^{\sqrt{z-3y^2}} dxdzdy\] Task: change the order of integral to dxdydz

hartnn (hartnn):

\( \large \int_0^1 \int_{3y^2}^{4-y^2}f(y,z)dzdy\) does this simplify things ? region between 2 parabolas...

OpenStudy (anonymous):

I tried, but failed :(

hartnn (hartnn):

plot z=y^2 z = 4-y^2 on same plane taking y as vertical and z as horizontal

hartnn (hartnn):

z=3y^2

OpenStudy (anonymous):

Hold on! I think I catch my mistake, just a minute!

hartnn (hartnn):

*staring at the screen*

OpenStudy (anonymous):